OpenStudy (anonymous):

Suppose you want to subtract: 9 - (-6). Re-write this subtraction as an ADDITION of signed numbers. 9 - (-6)

OpenStudy (anonymous):

9+6 because a minus a negative is the same as adding.

OpenStudy (anonymous):

what about -6-3?

OpenStudy (anonymous):

subtraction is defined by adding the opposite... a - b = a + (-b)

OpenStudy (anonymous):

no, -3 - 6 = -3 + (-6)

OpenStudy (anonymous):

yeah my bad, sorry got mixed up on that one.

OpenStudy (anonymous):

-7 - (-8)

OpenStudy (anonymous):

i dont understand these rewrite ones at all :(

OpenStudy (anonymous):

-7+8

OpenStudy (anonymous):

anymore?

OpenStudy (anonymous):

wanna answer my message?